OpenStudy (anonymous):

Simplify to lowest terms, if possible: (12 z^2)/(16 z^4)

OpenStudy (anonymous):

can you reduce the fraction \[\frac{12}{16}\]?

OpenStudy (anonymous):

what about the exponents? @satellite73

OpenStudy (anonymous):

we'll get there one step at a time

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

ok good now how about \[\frac{z^2}{z^4}\]?

OpenStudy (anonymous):

would you subtract them?

OpenStudy (anonymous):

yes you subtract the exponents, but since 4 is larger than 2, i would subtract with the z in the denominator

OpenStudy (anonymous):

\[\frac{12z^2}{16z^4}=\frac{3}{4z^{4-2}}=\frac{3}{4z^2}\]

OpenStudy (anonymous):

you are really just cancelling \[\frac{z^2}{z^4}=\frac{z\times z}{z\times z\times z\times z}=\frac{1}{z\times z}=\frac{1}{z^2}\]

OpenStudy (anonymous):

got it thanks!!

OpenStudy (anonymous):

yw